OpenStudy (anonymous):

f(x) = (x+5)^2 + (x-9)^2 how is this written in standard form... ?

OpenStudy (campbell_st):

I would expand and collect like terms f(x) = x^2 + 10x + 25 + x^2 - 18x + 81 thats the expansion

OpenStudy (anonymous):

so im getting 2x^2 - 8x +106

OpenStudy (campbell_st):

looks correct to me

OpenStudy (anonymous):

it doesnt work though.... yet i can tell that it is standard..

OpenStudy (campbell_st):

well make it monic.... divide by 2

OpenStudy (campbell_st):

or f(x) = 2(x^2 - 4x + 53)

OpenStudy (campbell_st):

standard for is\[ f(x) = a(x -h)^2 + k\] \[ = 2((x -2)^2 + 49)\] or \[ f(x) = 2(x -2)^2 + 98\]